# Backdoor Attack on Vertical Federated Graph Neural Network Learning

### v1

@@ 4.3.1 Impact of Multi-hop Trigger in BVG
4.3.2 Impact of Target Dataset Size
@@
4.3.3 Impact of Trigger Perturbation
@@
5 Conclusion
In this paper, we proposed the BVG method for backdoor attacks to Vertical Federated Graph Neural Network Learning (VFGNN). Utilizing a multi-hop trigger generation approach, the BVG method can perform efficient backdoor attacks with very limited knowledge of target class nodes. Extensive experiments using three GNN models across three datasets indicate that BVG achieves high attack success rates (ASR) with minimal impact on the main task accuracy (MTA). The evaluation of BVG efficacy under various defense methods highlights the robustness and efficiency of the proposed attack method, underscoring the necessity for advanced defense mechanisms in practical federated learning applications to counter such sophisticated backdoor attacks.

### v2

@@ B.1 Impact of Multi-hop Trigger in BVG
B.2 Impact of Backdoor Retention
To visually demonstrate the effect of Backdoor Retention, we conducted a total of nine experiments across three bottom models and three datasets, comparing the results with and without Backdoor Retention. In each experiment, we recorded the main task MTA, backdoor task ASR, and the current backdoor effect (effect) measured by the attacker after each epoch. The experimental results with GCN as the bottom model are shown in Figure 9, GAT results in Figure 10, and SGC results in Figure 11.
Figure 9: Impact of Backdoor Retention on Attack Stability Across Different Datasets. (GCN)
Figure 10: Impact of Backdoor Retention on Attack Stability Across Different Datasets. (GAT)
From the experiments without Backdoor Retention, we observed significant fluctuations in the backdoor task performance. This is mainly because the number of samples with known labels is too small, and their proportion is much lower compared to the thousands or tens of thousands of samples, making the trigger signal prone to being ”forgotten.” The evaluation of backdoor effect (effect) shows a strong correlation with the backdoor task performance, with its fluctuations almost perfectly reflecting the changes in the backdoor task.
In the experiments with Backdoor Retention, we observed a significant improvement in the stability of the backdoor task. This is particularly evident when GAT is used as the bottom model, where the backdoor performance fluctuations were almost completely eliminated. For GCN and SGC, the stability of the backdoor task also showed a clear improvement. This effect is primarily achieved because the attacker, through evaluating the backdoor effect, discards updates that negatively impact the backdoor effect and only accepts updates beneficial to the backdoor task, thereby ensuring the stability of the backdoor.
Figure 11: Impact of Backdoor Retention on Attack Stability Across Different Datasets. (SGC)
B.3 Impact of Target Dataset Size
@@
B.4 Impact of Trigger Perturbation
@@
Appendix C Attack Efficacy under Defense
Due to space constraints, we present the backdoor defense results for GAT and SGC as bottom models. Figures 14 and 15 show the defense results of the DP-SGD method, Figures 16 and 17 present the results of the GC method, Figures 18 and 19 illustrate the defense performance of the ANP method, and Figures 20 and 21 display the results of the ISO method.
The results indicate that different datasets and models respond similarly to these defense methods. DP-SGD and GC, which target label inference attacks, are almost ineffective against our attack. For the backdoor defense methods ANP and ISO, while the backdoor attack success rate decreases as model accuracy drops, these methods are still insufficient to effectively counter our attack.
Figure 14: The performance of BVG under DP-SGD defense (GAT)
Figure 15: The performance of BVG under DP-SGD defense (SGC)
Figure 16: The performance of BVG under GC defense (GAT)
Figure 17: The performance of BVG under GC defense (SGC)
Figure 18: The performance of BVG under ANP defense (GAT)
Figure 19: The performance of BVG under ANP defense (SGC)
Figure 20: The performance of BVG under ISO defense (GAT)
Figure 21: The performance of BVG under ISO defense (SGC)